# Steering Generative Models with Experimental Data for Protein Fitness Optimization

### v1

@@ 3 Problem setup
During inference, sequences can be sampled unconditionally from $p({\mathbf{x}})$, or sampling can be guided using a supervised model of the form $p({\mathbf{y}}|{\mathbf{x}})\propto\exp(f({\mathbf{x}})/\beta)$, where $f(\cdot)$ is a learned fitness predictor—also referred to as the classifier or value function. This predictor is trained on a small number of labeled sequence-fitness pairs (typically in the hundreds) to reflect practical data limitations. The goal of guided sampling is to generate protein sequences from the posterior distribution, $p({\mathbf{x}}|{\mathbf{y}})\propto p({\mathbf{x}})\exp(f({\mathbf{x}})/\beta)$. Details on training and guidance with the value function are provided in Section A.4 and Table A2. As an alternative steering method to guidance, we finetune the generative prior with labeled data using an autoregressive language model (ARLM) and DPO, which serves as a baseline. The strength of steering is tuned by method-specific hyperparameters.
@@ 4 Results
Table 2: Summary of datasets used in this work. Train and test fitness refer to the number of fitness labels used for training and testing the oracle. While the TrpB dataset has a lot more training labels, it may be more difficult to learn due to high amounts of epistatic effects between residues (non-additivity of mutation effects).
@@
389 117, 118, 119, 162, 166, 182, 183, 184, 185, 186, 227, 228, 230, 231, 301 N=15 $45.7e$ 4 75,618 23,313 Johnston et al. (2024)
CreiLOV Fluorescence 119 All N=119 $53.7e$ 5 6,842 2,401 Chen et al. (2023b)
We study two proteins, the TrpB enzyme (Johnston et al., 2024) and the CreiLOV fluorescent protein (Chen et al., 2023b) due to the availability of fitness data across many residues and the large number of homologous protein sequences found in their MSAs (Table 2). We focus protein fitness optimization to a design space of 15 residues in TrpB (only these positions are allowed to vary) and all 119 residues in CreiLOV. For each protein’s variants, we evaluate fitness by approximating it via a supervised oracle trained on a large amount of real data (Section A.2).
Based on the methods explained in Section A.1 and A.3, we trained generative priors on natural sequences from the MSA, focusing on continuous diffusion models (Continuous), discrete diffusion models with uniform (D3PM, UDLM) and absorbing state noising processes (MDLM), and autoregressive language models (ARLM) (Table 3). Overall, the trained models capture the natural distribution of protein sequences, with the D3PM models seeming to match the distribution the most closely while also generating sequences with high diversity (Fig. 4, Fig. A3). The two different diffusion models over continuous space show comparatively lower performance, and diffusing over the latent space of ESM embeddings does not boost performance on this task. The UDLM model has low performance due to mode collapse (Fig. A3, Fig. A4). Future work could finetune the pretrained diffusion protein language model (DPLM) as an MDLM (Wang et al., 2024b).
@@
D3PM 37.9 M finetuned from EvoDiff 38M-Uniform
@@
ARLM Language Model n/a 151 M autoregressive language model finetuned from ProGen2-small
Overall, we found that pretrained priors sample protein variants that have higher mean fitness, which corroborates previous studies finding that sequences with higher evolutionary likelihood are also likely to have higher fitness (Li et al., 2024a; Hie et al., 2023). Based on these results, we proceeded to perform remaining experiments with one model from each category of model type, namely the Continuous, D3PM, MDLM, and ARLM models.
Figure 4: Pretrained generative priors capture the target distribution of naturally occurring sequences that are homologous to TrpB (A-B) and CreiLOV (C-D), respectively. Lower perplexity corresponds to higher likelihood in the model. The diversity of sequences was computed as the average Shannon entropy of mutated positions with mean fitness corresponding to the oracle predictions. While the various models largely achieve comparable performance, the D3PM models capture the target distribution with the highest fidelity, whereas the UDLM model is prone to mode collapse. For each model, 1000 sequences were sampled and repeats were allowed to approximate the distribution. To approximate the target distribution, 1000 sequences were sampled from the MSA used for pretraining. Perplexity was calculated by passing generated sequences through the 764 M parameter ProGen2-base model. More details on model training can be found in Table 3 and Section A.3.
Figure 5: Pareto boundaries demonstrate the trade-off between generating sequences with high fitness and high diversity for TrpB (A-C) and CreiLOV (D-F). Sequences sampled from our generative models (Continuous, D3PM, and MDLM), after guidance with labeled fitness data, are enriched in high-fitness protein variants, and most methods show higher performance than the ARLM+DPO baseline. Larger circle indicates a stronger guidance strength hyperparameter, specified in Table A3. Each experiment was repeated using 10 different standardized sets of 200 unique sequences used for steering, each drawn from the D3PM prior, and error bars show standard deviation. Mean fitness and diversity were calculated based on 200 generated samples, with diversity calculated as the average Shannon entropy of amino acids at mutated positions. Unconditional refers to sequences sampled from the prior with no guidance.
@@ 4.2 Evaluating SGPO design choices
Impressively, steering with modest amounts of labeled data (200 sequence-fitness pairs) enables most models and methods to generate sequences with even higher fitness, while sacrificing some generation diversity (Fig. 5). In this low data regime, guidance with diffusion models outperforms DPO with language models, which works poorly on the CreiLOV dataset and only seems to steer generations well for a single hyperparameter on the TrpB dataset. CG and DAPS enable the strongest steerability overall, but DAPS outperforms CG for the continuous models (Fig. 5A, D). In general, guidance seems to work slightly better for uniform diffusion (D3PM) compared to absorbing state diffusion (MDLM). Overall, the continuous diffusion models do not perform as well as other models, as the prior does not capture the distribution of natural sequences with high fitness as well (Fig. 5A). NOS does not seem to allow for as much steerability, although we acknowledge that different hyperparameters could be tuned for NOS and DPO to potentially improve performance (Table A3). Finally, we conducted a closer analysis of the number of unique sequences generated by the steered models and confirmed that most models produce entirely novel sequences, suggesting that they are not over-steering (Fig. A5).
@@
Next, we performed adaptive optimization experiments, which mimic real-world protein engineering scenarios and follow a setup similar to batch Bayesian optimization: in each round, a batch of sequences is sampled, evaluated for fitness, and used to retrain a supervised value function that guides sampling from the pretrained prior. We focused on the D3PM and MDLM models with the CG and DAPS guidance strategies, as these combinations achieved the best performance in our earlier set of experiments (Fig. 5). Based on findings from these previous experiments, we selected the ideal guidance strength hyperparameter to balance fitness and diversity—ensuring high predicted fitness without significantly compromising sequence diversity (Table A3). For both guidance strategies, we employed an algorithm akin to Thompson sampling (Kandasamy et al., 2018; Russo et al., 2018), drawing a different value function from a frequentist ensemble of neural network regressors to guide the generation of each new sample (Yang et al., 2025). Pseudocode for our adaptive optimization algorithm is provided in Section A.5.
Figure 6: Maximum/mean fitness achieved improves over multiple iterations of steering in an adaptive setting similar to batch Bayesian optimization for TrpB (A-B) and CreiLOV (C-D). 100 sequences were sampled in each round. Within each round, an ensemble of 10 value functions (classifiers) was trained on fitness data from all previously queried samples, and each new sample was generated by the model guided with a value function sampled from the ensemble (akin to Thompson sampling). Only unique, novel samples were acquired. Guidance strength parameter is provided in Table A3. Error bars show standard deviation between 5 different random initializations.
This strategy outperforms baselines such as DPO with an ARLM or sampling just from the unconditional generative prior (Fig. 6), as sampled sequences achieve higher values of mean and maximum fitness. Furthermore, campaigns using an ensemble of value functions and “Thompson sampling” achieve higher maximum fitness than those using only a single value function for guidance (Table A4), which may be because these models enable more exploration of sequence space (Fig. A6). However, it is difficult to ascertain wither CG or DAPS works better as a guidance strategy, as the performance is highly dependent on the guidance strength hyperparameter, and the optimal hyperparameter will not typically be known in a real-world campaign. Because the oracle may not capture the true nature of the protein fitness landscape, we also suggest making relative comparisons here rather than absolute comparisons between model performance.
@@ 5 Discussion
In this work, we show that SGPO is an effective approach for protein fitness optimization, by capturing the distribution of natural protein sequences with a generative prior and then steering the generations with labeled data. We find that classifier guidance and posterior sampling (namely, DAPS) with discrete diffusion models have the highest performance, making plug-and-play guidance-based strategies more effective than finetuning language models; the latter can be difficult when only few fitness labels are available. Another major advantage of guidance is that only one hyperparameter (guidance strength) needs to be tuned, making it practical and user friendly for real-world applications. By contrast, for DPO, various hyperparameters need to be tuned, and the training process has to be monitored closely. Even for NOS, different parameters such as the step size, the number of steps, and the stability coefficient must be tuned together. A further advantage of guidance is the low computational cost required, as the prior model weights are not updated during guidance. Pretraining/finetuning to obtain each initial prior was achieved on a single H100 GPU in less than one hour while each individual guidance experiment took minutes; pretraining language models took several hours on a single GPU.
@@
There are certain limitations of our work. We focused on proteins with many homologous sequences and fitness as mostly native function, but it would be interesting to test SGPO on other protein fitness optimization tasks where the pretrained prior may not provide as much utility. We also focused on protein optimization where only $\approx 10^{2}$ fitness labels were available; different methods, such as RL, may perform better for applications where larger amounts of fitness data are available (Hie & Yang, 2022; Blalock et al., 2025). We focused on guidance strategies and did not test DPO or model finetuning-based methods with discrete diffusion models, but future work could adapt these methods for discrete diffusion (Borso et al., 2025). Furthermore, for TrpB and for language models, we manually mapped sequences back into the design space after generation (Section A.2), but explicitly building this into sampling techniques, such as inpainting in masked models (Blalock et al., 2025; Goel et al., 2024), may lead to improved performance. We did not consider insertions or deletions, but variable-length sequence generation could be considered in the future. Finally, we did not directly compare to existing approaches for protein engineering such as directed evolution for reasons explained in Section A.2.
@@
There are several promising directions for future work to improve and extend SGPO methods. For instance, we experimented with guiding generation using value functions sampled from a Gaussian process posterior, enabling principled Thompson sampling from a fully Bayesian perspective. However, the Gaussian process struggled to model high-dimensional protein representations, leading to poor performance. This limitation could potentially be addressed with better kernel choices (Wilson et al., 2015; Michael et al., 2024; Yang et al., 2025). Recent work has also begun to incorporate multi-objective optimization (Annadani et al., 2025; Tang et al., 2024; Chen et al., 2025) and uncertainty quantification (Wu et al., 2025) when guiding diffusion models. Simultaneously, alternatives to acquisition-function-based approaches are being developed to enable Bayesian optimization in large design spaces where enumeration is infeasible (Bal et al., 2024). Other emerging approaches—closer in spirit to flow matching—are being proposed for discrete data and may offer new opportunities for exploration (Davis et al., 2024; Stark et al., 2024; Tang et al., 2025). Finally, for masked diffusion models, strategies such as remasking could be explored to improve inference, particularly to enhance model amenability to guidance (Wang et al., 2025; Peng et al., 2025; Liu et al., 2024). It will also be interesting to further explore guidance in other discrete domains such as natural language and small molecules (Schiff et al., 2024).
@@ Acknowledgments
We would like to thank Hunter Nisonoff, Jacob Gershon, Lucas Arnoldt, Nate Gruver, and Chenghao Liu for helpful discussions and Francesca-Zhoufan Li for help with the TrpB dataset. We would also like to thank Nathaniel Blalock for guidance on how to use the CreiLOV dataset.
@@ A.1 Data for pretraining generative priors
The first step in our pipeline involves learning a generative prior on naturally occurring protein sequences to learn the distribution of those with high evolutionary likelihood. This prior is unconditional in the sense that no labeled fitness data is used for training. However, because we are optimizing protein variants for a desired fitness, we pretrained our generative prior on sequences homologous to the parent protein to be optimized (known as a multiple sequence alignment or MSA), either TrpB or CreiLOV. Likelihoods from MSAs have been captured by statistical models and have been shown to offer good zero-shot approximations of fitness. In other words, they capture mutational substitutions that are more favorable, based on the precedent of natural evolution.
We chose the TrpB (Johnston et al., 2024) and CreiLOV (Chen et al., 2023b) datasets due to the extensive number of sequences in their MSAs, which were obtained by running jackhmmer (Johnson et al., 2010) against Uniref90 for two iterations with the parent sequence of the fitness dataset as target. For the MSA, we only used sequences where the aligned portion was at least 75% the length of the parent sequence. For the diffusion model priors, we used the MSA that was aligned to the parent sequence, with gap tokens replaced by the corresponding amino acid found in the parent sequence, resulting in full, fixed-length pseudo-natural sequences. For the language model, the sequences used were the portions of the MSA that was aligned to the parent sequence, with gaps removed, thus resulting in variable length sequence fragments.

### v2

@@ 3 Problem setup
During inference, sequences can be sampled unconditionally from $p({\mathbf{x}})$, or sampling can be guided using a supervised model of the form $p({\mathbf{y}}|{\mathbf{x}})\propto\exp(f({\mathbf{x}})/\beta)$, where $f(\cdot)$ is a learned fitness predictor—also referred to as the classifier or value function. This predictor is trained on a small number of labeled sequence-fitness pairs (typically in the hundreds) to reflect practical data limitations. The goal of guided sampling is to generate protein sequences from the posterior distribution, $p({\mathbf{x}}|{\mathbf{y}})\propto p({\mathbf{x}})\exp(f({\mathbf{x}})/\beta)$. We use a computational oracle to acquire and evaluate fitness labels $\mathbf{y}$}, to simulate how fitness would be measured in a real-world campaign. Details on training and guidance with the value function are provided in Section A.4 and Table A2. As an alternative steering method to guidance, we finetune the generative prior with labeled data using an autoregressive language model (ARLM) and DPO, which serves as a baseline. We further compare to a baseline of latent space Bayesian optimization. The strength of steering is tuned by method-specific hyperparameters.
@@ 4 Results
Table 2: Summary of datasets used in this work. Train and test fitness refer to the number of fitness labels used for training and testing the oracle. We focus on TrpB and CreiLOV, with some of the GB1 results moved to the Appendix. While the TrpB dataset has a lot more training labels, it may be more difficult to learn due to relatively high amounts of epistatic effects between residues (non-additivity of mutation effects).
@@
389 117, 118, 119, 162, 166, 182, 183, 184, 185, 186, 227, 228, 230, 231, 301 N=15 $5.7e4$ 75,618 23,313 Johnston et al. (2024)
CreiLOV Fluorescence 119 All N=119 $3.7e5$ 6,842 2,401 Chen et al. (2023c)
GB1 Binding 56 All N=56 126126 3.9e6 9.6e4 Olson et al. (2014)
We study three proteins, the TrpB enzyme (Johnston et al., 2024), the CreiLOV fluorescent protein (Chen et al., 2023c), and the GB1 binding protein (Olson et al., 2014) due to the availability of fitness data across many residues (Table 2). We focus protein fitness optimization to a design space of 15 residues in TrpB (only these positions are allowed to vary) and all 119 and 56 residues in CreiLOV and GB1, respectively. For each protein’s variants, we evaluate fitness by approximating it via a supervised oracle trained on a large amount of real data (Section A.2).
@@
D3PM 37.9 M finetuned from EvoDiff 38M-Uniform (Alamdari et al., 2024)
@@
ARLM Language Model n/a 151 M autoregressive language model finetuned from ProGen2-small (Nijkamp et al., 2023)
Based on the methods explained in Section A.1 and A.3, we trained generative priors on natural sequences from the MSA, focusing on continuous diffusion models (Continuous), discrete diffusion models with uniform (D3PM, UDLM) and absorbing state noising processes (MDLM), and autoregressive language models (ARLM) (Table 3). Overall, the trained models capture the natural distribution of protein sequences, with the D3PM models seeming to match the distribution the most closely while also generating sequences with high diversity (Fig. 4, Fig. A3). The two different diffusion models over continuous space show comparatively lower performance, and diffusing over the latent space of ESM embeddings does not boost performance on this task. The UDLM model has low performance due to mode collapse (Fig. A3, Fig. A5). Future work could finetune the pretrained diffusion protein language model (DPLM) as an MDLM (Wang et al., 2024).
Overall, we found that pretrained priors sample protein variants that have higher mean fitness, which corroborates previous studies finding that sequences with higher evolutionary likelihood are also likely to have higher fitness (Li et al., 2025a; Hie et al., 2023). Based on these results, we proceeded to perform remaining experiments with one model from each category of model type, namely the Continuous, D3PM, MDLM, and ARLM models.
Figure 4: Pretrained generative priors capture the target distribution of naturally occurring sequences that are homologous to TrpB (A-B) and CreiLOV (C-D), respectively. Lower perplexity corresponds to higher likelihood in the model. The diversity of sequences was computed as the average Shannon entropy of mutated positions with mean fitness corresponding to the oracle predictions. While the various models largely achieve comparable performance, the D3PM models capture the target distribution with the highest fidelity, whereas the UDLM model is prone to mode collapse. For each model, 1000 sequences were sampled and repeats were allowed to approximate the distribution. To approximate the target distribution, 1000 sequences were sampled from the MSA used for pretraining. Perplexity was calculated by passing generated sequences through the 764 M parameter ProGen2-base model. More details on model training can be found in Table 3 and Section A.3, and GB1 results are provided in Fig. A4.
Figure 5: Pareto boundaries demonstrate the trade-off between generating sequences with high fitness and high diversity for TrpB (A-C), CreiLOV (D-F), and GB1 (G-I). Sequences sampled from the generative models (Continuous, D3PM, and MDLM), after guidance with labeled fitness data, are enriched in high-fitness protein variants, and most methods show higher performance than the ARLM+DPO baseline. Larger circle indicates a stronger guidance strength hyperparameter (excluding NOS), specified in Table A3. Each experiment was repeated using 10 different standardized sets of 200 unique sequences used for steering, each drawn from the D3PM prior, and error bars show standard deviation. Mean fitness and diversity were calculated based on 200 generated samples, with diversity calculated as the average Shannon entropy of amino acids at mutated positions. Unconditional refers to sequences sampled from the prior with no guidance.
@@ 4.2 Evaluating SGPO design choices
Impressively, steering with modest amounts of labeled data (200 sequence-fitness pairs) enables most models and methods to generate sequences with even higher fitness, while sacrificing some generation diversity (Fig. 5). In this low data regime, guidance with diffusion models outperforms DPO with language models; the latter does not enable as much steerability. CG and DAPS enable the strongest steerability overall, but DAPS outperforms CG for the continuous models (Fig. 5A, D). In general, guidance seems to work similarly for uniform diffusion (D3PM) and to absorbing state diffusion (MDLM). Overall, the continuous diffusion models do not perform as well as other models, as the prior does not capture the distribution of natural sequences with high fitness as well (Fig. 5A). NOS does not seem to allow for as much steerability, despite an extensive hyperparameter scan (Table A3). Finally, we conducted a closer analysis of the number of unique sequences generated by the steered models and confirmed that most models produce entirely novel sequences, suggesting that they are not over-steering (Fig. A6).
@@
Next, we performed adaptive optimization experiments, which mimic real-world protein engineering scenarios and follow a setup similar to batch Bayesian optimization: in each round, a batch of sequences is sampled, evaluated for fitness, and used to retrain a supervised value function that guides sampling from the pretrained prior. We focused on the MDLM models with the CG and DAPS guidance strategies, as these combinations achieved the best performance in our earlier set of experiments (Fig. 5). Based on findings from these previous experiments, we selected the ideal guidance strength hyperparameter to balance fitness and diversity—ensuring high predicted fitness without significantly compromising sequence diversity (Table A3). For both guidance strategies, we employed an algorithm akin to Thompson sampling (Kandasamy et al., 2018; Russo et al., 2018), drawing a different value function from a frequentist ensemble of neural network regressors to guide the generation of each new sample (Yang et al., 2025b). Pseudocode for our adaptive optimization algorithm is provided in Section A.5.
Plug-and-play guidance strategies outperform baselines such as DPO with an ARLM, sampling just from the unconditional generative prior, and latent space Bayesian optimization with APEXGo (Fig. 6): Sampled sequences achieve higher values of mean and maximum fitness. Furthermore, campaigns using an ensemble of value functions and “Thompson sampling” achieve higher maximum fitness than those using only a single value function for guidance (Table A4), which may be because these models enable more exploration of sequence space (Fig. A7). However, it is difficult to ascertain wither CG or DAPS works better as a guidance strategy, as the performance is highly dependent on the guidance strength hyperparameter, and the optimal hyperparameter will not typically be known in a real-world campaign. Because the oracle may not capture the true nature of the protein fitness landscape, we also suggest making relative comparisons here rather than absolute comparisons between model performance.
Figure 6: Maximum/mean fitness achieved improves over multiple iterations of steering in an adaptive setting similar to batch Bayesian optimization for TrpB (A), CreiLOV (B), and GB1 (C). 100 sequences were sampled in each round. Within each round, an ensemble of 10 value functions (classifiers) was trained on fitness data from all previously queried samples, and each new sample was generated by the MDLM model guided with a value function sampled from the ensemble (akin to Thompson sampling). Only unique, novel samples were acquired. Guidance strength parameter is provided in Table A3. Error bars show standard deviation between 5 different random initializations.
@@ 5 Discussion
In this work, we conduct a comprehensive study of SGPO methods and demonstrate that it is an effective approach for protein fitness optimization, by capturing the distribution of natural protein sequences with a generative prior and then steering the generations with labeled data. We find that DAPS with discrete diffusion models has the highest performance overall, and plug-and-play guidance-based strategies are generally more effective than finetuning language models; the latter can be difficult when only few fitness labels are available. SGPO approaches also outperform latent space Bayesian optimization (namely APEXGo), which we attribute to the difficulty in calibrating the trust region in very low-data regimes with limited rounds of optimization and the fact that latent space Bayesian optimization relies heavily on the structure of the latent space learned during generation model training, which can limit extrapolation to high-fitness but unnatural variants.
@@
Using plug-and-play guidance approaches has other advantages. First, only one hyperparameter (guidance strength) needs to be tuned. In real-world engineering scenarios, even in the absence of ground truth fitness labels, one practical approach to selecting the guidance strength is to scan over values and choose the highest setting for which $n$ generated sequences remain unique and novel relative to previously measured sequences, where $n$ corresponds to the screening throughput available for the next round. By contrast, for DPO, various hyperparameters need to be tuned, and the training process has to be monitored closely. Even for NOS, different parameters such as the step size, the number of steps, and the stability coefficient must be tuned together. A further advantage of guidance is the low computational cost required, as the prior model weights are not updated during guidance. Pretraining/finetuning to obtain each initial prior was achieved on a single H100 GPU in less than one hour while each individual guidance experiment took minutes; pretraining language models took several hours on a single GPU.
@@
Still, there are certain limitations of our work. We focused on proteins with fitness as mostly native function, but it would be interesting to test SGPO on other protein fitness optimization tasks where the pretrained prior may not provide as much utility. We also focused on protein optimization where only $\approx 10^{2}$ fitness labels were available; different methods, such as RL, may perform better for applications where larger amounts of fitness data are available (Hie & Yang, 2022; Blalock et al., 2025). We focused on guidance strategies and did not test DPO or model finetuning-based methods with discrete diffusion models, but future work could adapt these methods for discrete diffusion (Borso et al., 2025). Furthermore, for TrpB and for language models, we manually mapped sequences back into the design space after generation (Section A.2), but explicitly building this into sampling techniques, such as inpainting in masked models (Blalock et al., 2025; Goel et al., 2025) may lead to improved performance. We did not consider insertions or deletions, but variable-length sequence generation could be considered in the future. Finally, we did not directly compare to existing approaches for protein engineering such as directed evolution for reasons explained in Section A.2.
@@
There are several promising directions for future work to improve and extend SGPO methods. For instance, we experimented with guiding generation using value functions sampled from a Gaussian process posterior, enabling principled Thompson sampling from a fully Bayesian perspective. However, the Gaussian process struggled to model high-dimensional protein representations, leading to poor performance. This limitation could potentially be addressed with better kernel choices (Wilson et al., 2016; Michael et al., 2024; Yang et al., 2025b). Recent work has also begun to incorporate multi-objective optimization (Annadani et al., 2025; Tang et al., 2025a; Li et al., 2024b; Chen et al., 2025) and uncertainty quantification (Wu et al., 2025) when guiding diffusion models. Simultaneously, alternatives to acquisition-function-based approaches are being developed to enable Bayesian optimization in large design spaces where enumeration is infeasible (Bal et al., 2025). Other emerging approaches—closer in spirit to flow matching—are being proposed for discrete data and may offer new opportunities for exploration (Davis et al., 2024; Stark et al., 2024; Tang et al., 2025b). Finally, for masked diffusion models, strategies such as remasking or scheduling could be explored to improve inference, particularly to enhance model amenability to guidance (Wang et al., 2025b; Peng et al., 2025; Liu et al., 2025; Amin et al., 2025a). It will also be interesting to further explore guidance in other discrete domains such as natural language and small molecules (Schiff et al., 2024).
@@ Acknowledgments
This work was supported by a U.S. Army Research Office cooperative agreement (W911NF-19-2-0026 to F.H.A.) and an Amgen Chem-Bio Engineering award. J.Y. is also supported by the NSF Graduate Research Fellowship Program and the Google PhD Fellowship. We would like to thank Hunter Nisonoff, Jacob Gershon, Lucas Arnoldt, and Chenghao Liu for helpful discussions and Francesca-Zhoufan Li for help with the TrpB dataset. We would also like to thank Nate Gruver for help with the NOS implementation, Filippo Stocco for help with the DPO implementation, and Nathaniel Blalock for guidance on how to use the CreiLOV dataset.
@@ A.1 Data for pretraining generative priors
The first step in our pipeline involves learning a generative prior on naturally occurring protein sequences to capture the distribution of those with high evolutionary likelihood. This prior is unconditional in the sense that no labeled fitness data is used for training. However, because we are optimizing protein variants for a desired fitness, we pretrained our generative prior on sequences homologous to the parent protein to be optimized (known as a multiple sequence alignment or MSA): TrpB, CreiLOV, or GB1. Likelihoods from MSAs have been captured by statistical models and have been shown to offer good zero-shot approximations of fitness. In other words, they capture mutational substitutions that are more favorable, based on the precedent of natural evolution.
We focused on the TrpB (Johnston et al., 2024) and CreiLOV (Chen et al., 2023c) datasets due to the extensive number of sequences in their MSAs and compared to GB1 (Olson et al., 2014), which has comparatively fewer sequences. MSAs were obtained by running jackhmmer (Johnson et al., 2010) against Uniref90 for two iterations with the parent sequence of the fitness dataset as target. For the MSA, we only used sequences where the aligned portion was at least 75% the length of the parent sequence. We used the MSA that was aligned to the parent sequence, with gap tokens replaced by the corresponding amino acid found in the parent sequence, resulting in full, fixed-length pseudo-natural sequences. For GB1, we augmented the training set with synthetic data, namely all proteins with a single mutation to sequences in the MSA. For the language models on TrpB and CreiLOV, some sequences were randomly mutated by a single position near the beginning of the sequence, to prevent mode collapse during autoregressive generation.